OpenStudy (anonymous):

x^2-5x+6=0 Solve by completing the square

OpenStudy (anonymous):

x^2-5x+6 x^2-3x-2x+6 x(x-3)-2(x-3) (x-2)(x-3)

OpenStudy (anonymous):

can you explain how you got the -3x and -2x in the second step?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

standard form\[ax^2 + bx + c\] now i split b into y and z, so that \[y + z = b\] \[y \times z = ac\] in this case, b is -5 and ac is 6 so i chose y = -3 and z = -2 -3 + -2 = -5 -3 * -2 = 6

OpenStudy (anonymous):

with y and z satisfy these conditions, you will be able to factor nicely

OpenStudy (anonymous):

hmm so i didnt really need to do the whole (b/2)^2 thing cuz it just factored from the start?

OpenStudy (anonymous):

ohhhhhh wooops completing the squareee, the method i used is called factoring

OpenStudy (anonymous):

lemme try the other way

OpenStudy (anonymous):

ya but you got the right answer that way haha

OpenStudy (anonymous):

i like your way better

OpenStudy (anonymous):

x^2-5x+6 x^2-5x = -6 \[x^2 -5x + \frac{25}{4} = -6 + \frac{25}{4}\] \[(x-\frac{5}{2})^2=\frac{1}{4}\] \[x-\frac{5}{2}=+-1/2\] x = 6/2 = 3 x = 4/2 = 2

OpenStudy (anonymous):

okay i get it now thank you

OpenStudy (anonymous):

you re welcome